# Estimating Causal Effects for Binary Outcomes Using Per-Decision Inverse Probability Weighting

### v2

@@ 3 An Estimator Using Per-Decision Inverse Probability Weights
(Small-sample correction for inference.) The sandwich estimator for the variance in Theorem 1 can be anti-conservative with small sample size. To correct this, we adopt two small-sample correction techniques: (i) Similar to the correction to the sandwich variance estimator by [29], we pre-multiply the vector of each individual’s residual, $\epsilon_{t}(\hat{\alpha},\hat{\beta})$, by the inverse of the identity matrix minus the leverage for that individual; (ii) Similar to the correction to the critical value for inference by [6], we use as the critical value the quantile of a $t$-distribution, $t^{-1}_{n-p-q}(1-\eta/2)$, where $\eta$ is the significance level and $p$ and $q$ are the dimensions of $\beta$ and $\alpha$, respectively.
@@ 4 Simulation
4.1 Overview
We studied the finite-sample performance of our proposed estimator and three competitors: the EMEE by [11], and two generalized estimating equations [30]. We will refer to our estimator as “pd-EMEE”, where “pd” stands for per-decision.
@@ 4.2 Generative Model
We set $I_{t}\equiv 1$ and $T=100$. The generative model depends on two parameters, $\Delta$ (length of the proximal outcome window) and $p_{a}$ (constant randomization probability). $A_{t}\in\{0,1\}$ was generated from $\text{Bernoulli}(p_{a})$. A scalar time-varying covariate, $Z_{t}\in\{0,1,2\}$, was generated independently from all previous variables (i.e., $Z_{t}\perp\{Z_{s},A_{s},R_{s+1}:1\leq s\leq t-1\}$) with
@@
In Supplementary Material F we showed that for $1\leq t\leq T-\Delta$,
$E(Y_{t,\Delta}(\bar{A}_{t-1},a_{t},\bar{0}_{\Delta-1})\mid H_{t}(\bar{A}_{t-1}),I_{t}(\bar{A}_{t-1})=1)=\{1-0.5^{(\Delta+0.5-0.5Z_{t})/\Delta}\cdot(3/C)^{\Delta-1}\}\cdot e^{a_{t}(0.1+0.2Z_{t})}.$ (10)
@@
There is substantial variability in $R_{t+1}$ and $Y_{t,\Delta}$. When we vary $\Delta$ from 2 to 10 and vary $Z_{t}\in\{0,1,2\}$, (8) varies from 0.59 to 0.97, (9) varies from 0.38 to 0.80, and (10) varies from 0.41 to 0.58 when setting $a_{t}=0$ and from 0.57 to 0.80 when setting $a_{t}=1$.
We constructed this generative model based on three considerations. First, in order to fairly assess how the efficiency improvement is affected by $\Delta$ and $p_{a}$, we needed $\text{CEE}(Z_{t})$ to not depend on $\Delta$ or $p_{a}$. Thus, the distribution of the sub-outcome $R_{t}$ would depend on $\Delta$ because there are $\Delta$ sub-outcomes in $Y_{t,\Delta}$. Second, we wanted to construct a complicated $E(Y_{t,\Delta}(\bar{A}_{t-1},0,\bar{0}_{\Delta-1})\mid H_{t},I_{t}=1)$ to demonstrate the robustness of the estimator against misspecified working models. Third, we wanted $Z_{t}$ to depend on $\Delta$ as little as possible. In a real application, $Z_{t}$ (a covariate) would not depend on $\Delta$ (a choice in the analysis). However, we found it difficult to achieve while keeping the form of $\text{CEE}(Z_{t})$ simple. The current generating distribution of $Z_{t}$ (7) is the least dependent on $\Delta$ among the ones we came up with, and the three probabilities in (7) are all very close to 1/3 for $\Delta\geq 2$.
@@ 4.3 Simulation Result
We first conducted simulations with different $\Delta$: $\Delta=3$ and $\Delta=10$. We used a constant randomization probability $p_{a}=0.2$. We considered two estimands with different $S_{t}$: the first estimand is $\text{CEE}(S_{t})=\beta_{0}$ with $S_{t}=1$, and the second estimand is $\text{CEE}(S_{t})=\beta_{1}+\beta_{2}Z_{t}$ with $S_{t}=(1,Z_{t})$. They represent common primary and secondary analyses of MRTs: $\beta_{0}$ is the fully marginal causal excursion effect, and $\beta_{1}+\beta_{2}Z_{t}$ is the effect moderated by $Z_{t}$.
@@
Four estimators were compared: our proposed pd-EMEE, the EMEE by [11], the log-linear generalized estimating equations with independent correlation structure (GEE.ind), and the log-linear generalized estimating equations with exchangeable correlation structure (GEE.exch). For pd-EMEE and EMEE, we used the correct model of CEE for both estimands and the incorrect working model $\exp\{g(H_{t})^{T}\alpha\}=\exp(\alpha_{0}+\alpha_{1}Z_{t})$ for $E(Y_{t,\Delta}(\bar{A}_{t-1},0,\bar{0}_{\Delta-1})\mid H_{t},I_{t}=1)$. For GEE.ind and GEE.exch, we use the marginal mean model $\exp(\alpha_{0}+\alpha_{1}Z_{t}+\beta_{0}A_{t})$ for the first estimand and $\exp\{\alpha_{0}+\alpha_{1}Z_{t}+A_{t}(\beta_{1}+\beta_{2}Z_{t})\}$ for the second estimand. The marginal mean models used by the GEEs are misspecified because the part involving the $\alpha$’s are incorrect. See Supplementary Material G for the exact form of the log-linear GEEs used.
@@
When $\Delta=3$, the true parameter values are $\beta_{0}^{*}=0.283$ (obtained numerically by plugging (10) into (2)), $\beta_{1}^{*}=0.1$, and $\beta_{2}^{*}=0.2$. Both pd-EMEE and EMEE consistently estimate all parameters of interest with close to nominal confidence interval coverage (top halves of Tables 2 and 3). The confidence interval coverage is improved by the small sample correction (Remark 2). pd-EMEE is slightly more efficient than EMEE: for example, when $n=100$, the relative efficiency between pd-EMEE and EMEE for estimating $\beta_{0}$, $\beta_{1}$, and $\beta_{2}$ is $1.080.026^{2}/0.025^{2}=$ 1.08, $1.120.037^{2}/0.035^{2}=$ 1.12, and $1.150.029^{2}/0.027^{2}=$ 1.15, respectively. These values roughly translate to 7% ($=1-1/1.08$), 11%, and 13% savings in sample size if one were to power an MRT using pd-EMEE instead of EMEE. Both GEE.ind and GEE.exch are inconsistent because the marginal mean model is misspecified. The performance of $\hat{\alpha}_{0}$ and $\hat{\alpha}_{1}$ is not presented because $\alpha_{0}$ and $\alpha_{1}$ are nuisance parameters and thus not of interest.
@@
When $\Delta=10$, the true parameter values are $\beta_{0}^{*}=0.304$, $\beta_{1}^{*}=0.1$, and $\beta_{2}^{*}=0.2$. The conclusions regarding consistency and confidence interval coverage are the same as when $\Delta=3$ (bottom halves of Tables 2 and 3). Here pd-EMEE is substantially more efficient than EMEE: for example, when $n=100$, the relative efficiency between pd-EMEE and EMEE for estimating $\beta_{0}$, $\beta_{1}$, and $\beta_{2}$ is $1.450.065^{2}/0.054^{2}=$ 1.45, $1.390.092^{2}/0.078^{2}=$ 1.39, and $1.400.064^{2}/0.054^{2}=$ 1.40, respectively, which roughly translates to 31%, 28%, and 29% savings in sample size. The efficiency improvement is greater compared to when $\Delta=3$.
@@
Table 2: Performance of pd-EMEE, EMEE, GEE.ind and GEE.exch when estimating $\beta_{0}$ ($\Delta$ = 3 and 10). Based on 1,000 simulation replications.
@@
  • pd-EMEE, our proposed estimator; EMEE, the estimator in [11]; GEE.ind, generalized estimating equations with independent working correlation structure; GEE.exch, generalized estimating equations with exchangeable working correlation structure; SD, standard deviation; RMSE, root mean squared error; CP, $95\%$ confidence interval coverage probability before (unadj) and after (adj) small-sample correction.
Table 3: Performance of pd-EMEE, EMEE, GEE.ind and GEE.exch when estimating $(\beta_{1},\beta_{2})$ ($\Delta$ = 3 and 10). Based on 1,000 simulation replications.
@@
  • pd-EMEE, our proposed estimator; EMEE, the estimator in [11]; GEE.ind, generalized estimating equations with independent working correlation structure; GEE.exch, generalized estimating equations with exchangeable working correlation structure. SD, standard deviation; RMSE, root mean squared error; CP, $95\%$ confidence interval coverage probability before (unadj) and after (adj) small-sample correction.
We then investigated how the relative efficiency between pd-EMEE and EMEE depends on $\Delta$ (proximal outcome window length) and $p_{a}$ (randomization probability) by varying $\Delta$ and $p_{a}$ separately. When $\Delta$ varies from 1 to 10 with $p_{a}=0.2$, the relative efficiency increases from 1 to 1.45 (left panel of Figure 2). When $p_{a}$ varies from 0.1 to 0.8 with $\Delta=3$, the relative efficiency increases from 1.05 to 1.76 (right panel of Figure 2). Therefore, a larger $\Delta$ or a larger $p_{a}$ makes the proposed pd-EMEE method even more efficient than the EMEE method. This phenomenon was expected and explained in Section 3.
Figure 2: Relative efficiency $\text{Var}(\hat{\beta}^{\text{EMEE}})/\text{Var}(\hat{\beta}^{\text{pd-EMEE}})$ against $\Delta$ (proximal outcome window length) and $p$ (randomization probability).

### v3

@@ 2.4 Identification of Parameters
(Small-sample correction for inference.) The sandwich estimator for the variance in Theorem 1 can be anti-conservative with a small sample size. To correct this, we adopt two small-sample correction techniques: (i) Similar to the correction to the sandwich variance estimator by Mancl and DeRouen (2001), we pre-multiply the vector of each individual’s residual, $\epsilon_{t}(\hat{\alpha},\hat{\beta})$, by the inverse of the identity matrix minus the leverage for that individual; (ii) Similar to the correction to the critical value for inference by Liao and others (2016), we use as the critical value the quantile of a $t$-distribution, $t^{-1}_{n-p-q}(1-\eta/2)$, where $\eta$ is the significance level and $p$ and $q$ are the dimensions of $\beta$ and $\alpha$, respectively.
4 pd-EMEE2: Further Improving Efficiency Using Projection
Following suggestions by a reviewer, we developed a variation of the pd-EMEE $\hat{\beta}$ using the projection idea from semiparametric efficiency theory. We refer to this variation as pd-EMEE2 and denote it $\tilde{\beta}$. In particular, let $\xi_{i}(\beta)=\sum_{t=1}^{T}I_{it}e^{-A_{it}S_{it}^{T}\beta}Y_{it,\Delta}M_{it}W_{it}\{A_{it}-\tilde{p}_{t}(S_{it})\}S_{it}$ denote the bottom portion of $U_{i}(\alpha,\beta)$ in (6) with $\epsilon_{it}(\alpha,\beta)$ replaced by $Y_{it,\Delta}$. Equation (4) implies the moment condition $E\{\xi_{i}(\beta^{*})\}=0$ (proven in Lemma 7 in the Supplementary Material). The efficiency of the estimating function $\xi_{i}(\beta)$ can be improved by subtracting from $\xi_{i}(\beta)$ its projection on the score functions for the treatment assignment probability (Bickel and others, 1993; Murphy and others, 2001; Shi and Dempsey, 2023). The projection is given by $\sum_{u=1}^{T}[E\{\xi_{i}(\beta)\mid H_{iu},A_{iu}\}-E\{\xi_{i}(\beta)\mid H_{iu}\}]$ (Robins, 1999). We show in Supplementary Material G that under an additional working assumption (Assumption 8 in Supplementary Material G), we have
$~{}~{}~{}~{}\xi_{i}(\beta)-\sum_{u=1}^{T}[E\{\xi_{i}(\beta)\mid H_{iu},A_{iu}\}-E\{\xi_{i}(\beta)\mid H_{iu}\}]$
$=\sum_{t=1}^{T}I_{it}e^{-A_{it}S_{it}^{T}\beta}M_{it}\{A_{it}-\tilde{p}_{t}(S_{it})\}S_{it}\bigg{[}Y_{it,\Delta}W_{it}-E(Y_{it,\Delta}W_{it}|H_{it},A_{it})$
$~{}~{}~{}~{}-\sum_{u=t+1}^{t+\Delta-1}\left\{E(Y_{it,\Delta}W_{it}|H_{iu},A_{iu})-E(Y_{it,\Delta}W_{it}|H_{iu})\right\}\bigg{]}$
$+\sum_{t=1}^{T}I_{it}\tilde{p}_{t}(S_{it})\{1-\tilde{p}_{t}(S_{it})\}S_{it}\left\{e^{-S_{it}^{T}\beta}E(Y_{it,\Delta}W_{it}|H_{it},A_{it}=1)-E(Y_{it,\Delta}W_{it}|H_{it},A_{it}=0)\right\}.$ (7)
This motivates the following estimating function
$\tilde{U}_{i}(\beta,\mu)$ $=\sum_{t=1}^{T}I_{it}e^{-A_{it}S_{it}^{T}\beta}M_{it}\{A_{it}-\tilde{p}_{t}(S_{it})\}S_{it}\bigg{[}Y_{it,\Delta}W_{it}-\mu_{0}(H_{it},A_{it})$
$~{}~{}~{}~{}-\sum_{u=t+1}^{t+\Delta-1}\left\{\mu_{u-t}(H_{iu},A_{iu})-p_{u}(H_{iu})\mu_{u-t}(H_{iu},1)-(1-p_{u}(H_{iu}))\mu_{u-t}(H_{iu},0)\right\}\bigg{]}$
$+\sum_{t=1}^{T}I_{it}\tilde{p}_{t}(S_{it})\{1-\tilde{p}_{t}(S_{it})\}S_{it}\left\{e^{-S_{it}^{T}\beta}\mu_{0}(H_{it},1)-\mu_{0}(H_{it},0)\right\},$ (8)
where $\mu=(\mu_{0},\mu_{1},\ldots,\mu_{\Delta-1})$ is a collection of nuisance parameters with $\mu_{s}(H_{iu},A_{iu})$ denoting a working model for $E(Y_{i,u-s,\Delta}W_{i,u-s}|H_{iu},A_{iu})$. We propose to compute the estimator $\tilde{\beta}$ based on the estimating function $\tilde{U}$ using a two-step approach depicted in Algorithm 1. The asymptotic property of $\tilde{\beta}$ is described in Theorem 2, with proof in Supplementary Material G.2.
Algorithm 1 The pd-EMEE2 estimator $\tilde{\beta}$
Step 1: For each $s=0,1,\ldots,\Delta-1$, fit $E(Y_{i,u-s,\Delta}W_{i,u-s}|H_{iu},A_{iu})$ by pooling across all $1\leq u\leq T$ and denote the fitted models by $\hat{\mu}_{s}(H_{iu},A_{iu})$. Denote by $\hat{\mu}$ the collection of $\hat{\mu}_{s}$ for $10\leq s\leq\Delta-$ 1.
Step 2: Obtain $\tilde{\beta}$ by solving $\frac{1}{n}\sum_{i=1}^{n}\tilde{U}_{i}(\beta;\hat{\mu})=0$ with $\tilde{U}_{i}(\beta,\mu)$ defined in (8).
Theorem 2
Assume the causal effect model (5) and Assumptions 1-4 in Section 2.4. Let $\beta^{*}$ denote the true value of $\beta$ corresponding to the data generating distribution $P_{0}$. Define $\partial_{\beta}\tilde{U}_{i}(\beta;\hat{\mu})=\partial\tilde{U}_{i}(\beta;\hat{\mu})/\partial\beta^{T}$. Under regularity conditions, $\tilde{\beta}$ is consistent for $\beta^{*}$ and $\sqrt{n}(\tilde{\beta}-\beta^{*})$ is asymptotically normal with mean zero and variance-covariance matrix $\Sigma$, and $\Sigma$ can be consistently estimated by the lower block diagonal $(p\times p)$ entry of the matrix
$\bigg{\{}\frac{1}{n}\sum_{i=1}^{n}\partial_{\beta}\tilde{U}_{i}(\tilde{\beta};\hat{\mu})\bigg{\}}^{-1}\bigg{\{}\frac{1}{n}\sum_{i=1}^{n}\tilde{U}_{i}(\tilde{\beta};\hat{\mu})\tilde{U}_{i}(\tilde{\beta};\hat{\mu})^{T}\bigg{\}}\bigg{\{}\big{[}\frac{1}{n}\sum_{i=1}^{n}\partial_{\beta}\tilde{U}_{i}(\tilde{\beta};\hat{\mu})\big{]}^{-1}\bigg{\}}^{T}.$
Remark 3
(Robustness and efficiency improvement.) Similar to Remark 1, the consistency of $\tilde{\beta}$ is robust against misspecified regression models $\hat{\mu}$. Equation (7) is rooted in the semiparametric efficiency theory, and one can view (7) as identifying a collection of working models that can potentially lead to more efficient $\beta$ estimators than the working model $g(H_{it})^{T}\alpha$ in (6). In other words, the pd-EMEE2 $\tilde{\beta}$ can be more efficient than the pd-EMEE $\hat{\beta}$. Note that Assumption 8 was made to facilitate a closed-form (7), and this working assumption is not required by Theorem 2. Furthermore, the efficiency gain by $\tilde{\beta}$ does not require this working assumption, as is supported by numerical evidence (Sections 5 and 6).
Remark 4
(Flexible modeling choice for $\mu$.) The choice of models for the conditional expectations $E(Y_{i,u-s,\Delta}W_{i,u-s}|H_{iu},A_{iu})$ in Algorithm 1 can be very flexible, and one can take advantage of the maximum property to further improve efficiency. For example, one can use Poisson regression or logistic regression with rescaled $Y_{i,u-s,\Delta}W_{i,u-s}$ that are within $[0,1]$ (and scale them back after the logistic regression fit). One can also include functions and spline basis of the decision point index, $u$, in the model. In addition, the $(H_{iu},A_{iu})$-measurable part of $W_{i,u-s}$, i.e., $\prod_{j=u-s+1}^{u}\left\{\frac{\mathbbm{1}(A_{ij}=0)}{1-p_{j}(H_{ij})}\right\}^{\mathbbm{1}(\max_{u-s+1\leq v\leq j}R_{i,v}=0)}$, can be taken outside of the conditional expectation $E(Y_{i,u-s,\Delta}W_{i,u-s}|H_{iu},A_{iu})$, before a model is fitted to the conditional expectation, and doing this has the potential to improve efficiency. Furthermore, due to the maximum property, $Y_{i,u-s,\Delta}$ will be deterministically 1 if one of its sub-outcomes $R$ is included in $H_{iu}$ and equals 1, and taking advantage of this fact can further improve efficiency. Finally, due to the estimating function $\tilde{U}_{i}(\beta,\mu)$ being globally robust (i.e., it has expectation 0 at the true $\beta^{*}$ for any $\mu$; see Supplementary Material G.2), the nuisance parameter estimators $\hat{\mu}$ can converge at an arbitrarily slow rate and the asymptotic variance of $\tilde{\beta}$ does not need to account for the fact that $\mu$ is estimated (Cheng and others, 2023). This means that a wide range of machine learning and nonparametric models can be used to fit $\mu$.
@@ 5.1 Generative Model
We set $I_{t}\equiv 1$ and $T=100$. The generative model depends on two parameters, $\Delta$ (length of the proximal outcome window) and $p_{a}$ (constant randomization probability). $A_{t}\in\{0,1\}$ was generated from $\text{Bernoulli}(p_{a})$. A scalar time-varying covariate, $Z_{t}\in\{0,1,2\}$, was generated independently of all previous variables (i.e., $Z_{t}\perp\{Z_{s},A_{s},R_{s+1}:1\leq s\leq t-1\}$) with
@@
In Supplementary Material H we showed that for $1\leq t\leq T-\Delta$,
$E\{Y_{t,\Delta}(\bar{A}_{t-1},a_{t},\bar{0}_{\Delta-1})\mid H_{t}(\bar{A}_{t-1}),I_{t}(\bar{A}_{t-1})=1\}=\{1-0.5^{(\Delta+0.5-0.5Z_{t})/\Delta}\cdot(3/C)^{\Delta-1}\}\cdot e^{a_{t}(0.1+0.2Z_{t})}.$ (12)
@@
There is substantial variability in $R_{t+1}$ and $Y_{t,\Delta}$. When we vary $\Delta$ from 2 to 10 and vary $Z_{t}\in\{0,1,2\}$, (H.2) varies from 0.59 to 0.97, (H.3) varies from 0.38 to 0.80, and (12) varies from 0.41 to 0.58 when setting $a_{t}=0$ and from 0.57 to 0.80 when setting $a_{t}=1$.
We constructed this generative model based on three considerations. First, in order to fairly assess how the efficiency improvement is affected by $\Delta$ and $p_{a}$, we needed $\text{CEE}(Z_{t})$ to not depend on $\Delta$ or $p_{a}$. Thus, the distribution of the sub-outcome $R_{t}$ would depend on $\Delta$ because there are $\Delta$ sub-outcomes in $Y_{t,\Delta}$. Second, we wanted to construct a complicated $E\{Y_{t,\Delta}(\bar{A}_{t-1},0,\bar{0}_{\Delta-1})\mid H_{t},I_{t}=1\}$ to demonstrate the robustness of the estimator against misspecified working models. Third, we wanted $Z_{t}$ to depend on $\Delta$ as little as possible. In a real application, $Z_{t}$ (a covariate) would not depend on $\Delta$ (a choice in the analysis). However, we found it difficult to achieve while keeping the form of $\text{CEE}(Z_{t})$ simple. The current generating distribution of $Z_{t}$ (H.1) is the least dependent on $\Delta$ that we came up with, and the three probabilities in (H.1) are all very close to 1/3 for $\Delta\geq 2$.
@@ 5.2 Simulation Result
We studied the finite-sample performance of the proposed pd-EMEE, pd-EMEE2, and three competitors: the EMEE by Qian and others (2021), the log-linear generalized estimating equations with independent correlation structure (GEE.ind), and the log-linear generalized estimating equations with exchangeable correlation structure (GEE.exch) (Liang and Zeger, 1986).
@@
We considered two estimands with different $S_{t}$: the first estimand is $\text{CEE}(S_{t})=\beta_{0}$ with $S_{t}=1$, and the second estimand is $\text{CEE}(S_{t})=\beta_{1}+\beta_{2}Z_{t}$ with $S_{t}=(1,Z_{t})$. They represent common primary and secondary analyses of MRTs: $\beta_{0}$ is the fully marginal causal excursion effect, and $\beta_{1}+\beta_{2}Z_{t}$ is the effect moderated by $Z_{t}$. For pd-EMEE and EMEE, we used the correct model of CEE for both estimands and the incorrect working model $\exp\{g(H_{t})^{T}\alpha\}=\exp(\alpha_{0}+\alpha_{1}Z_{t})$ for $E(Y_{t,\Delta}(\bar{A}_{t-1},0,\bar{0}_{\Delta-1})\mid H_{t},I_{t}=1)$. For pd-EMEE2, we used the correct model of CEE and included only $(Z_{iu},A_{iu})$ in a linear regression fit for $\hat{\mu}_{s}(H_{iu},A_{iu})$ and thus the latter is misspecified. For GEE.ind and GEE.exch, we use the marginal mean model $\exp(\alpha_{0}+\alpha_{1}Z_{t}+\beta_{0}A_{t})$ for the first estimand and $\exp\{\alpha_{0}+\alpha_{1}Z_{t}+A_{t}(\beta_{1}+\beta_{2}Z_{t})\}$ for the second estimand. The marginal mean models used by the GEEs are misspecified because the part involving the $\alpha$’s is incorrect. See Supplementary Material I for the exact form of the log-linear GEEs used.
@@
We first conducted simulations with two different $\Delta$: $\Delta=3$ and $\Delta=10$. We used a constant randomization probability $p_{a}=0.2$. Results for $\Delta=3$ are shown in top halves of Tables 2 and 3, where the true parameter values are $\beta_{0}^{*}=0.283$ (obtained numerically by plugging (12) into (2)), $\beta_{1}^{*}=0.1$, and $\beta_{2}^{*}=0.2$. pd-EMEE, pd-EMEE2, and EMEE consistently estimate all parameters of interest with close to nominal confidence interval coverage. The confidence interval coverage is improved by the small sample correction (Remark 2). pd-EMEE is more efficient than EMEE: for example, when $n=100$, the relative efficiency between pd-EMEE and EMEE for estimating $\beta_{0}$, $\beta_{1}$, and $\beta_{2}$ is $1.080.026^{2}/0.025^{2}=$ 1.08, $1.120.037^{2}/0.035^{2}=$ 1.12, and $1.150.029^{2}/0.027^{2}=$ 1.15, respectively. These values roughly translate to 7% ($=1-1/1.08$), 11%, and 13% savings in sample size if one were to power an MRT using pd-EMEE instead of EMEE. Comparing pd-EMEE and pd-EMEE2, pd-EMEE2 is as or more efficient across $n=30,50\text{ or }100$, with the highest relative efficiency (against pd-EMEE) being $1.08$. Both GEE.ind and GEE.exch are inconsistent because the marginal mean model is misspecified. The performance of $\hat{\alpha}_{0}$ and $\hat{\alpha}_{1}$ is not presented because $\alpha_{0}$ and $\alpha_{1}$ are nuisance parameters and thus not of interest.
@@
Results for $\Delta=10$ are shown in bottom halves of Tables 2 and 3, where the true parameter values are $\beta_{0}^{*}=0.304$, $\beta_{1}^{*}=0.1$, and $\beta_{2}^{*}=0.2$. The conclusions regarding consistency and confidence interval coverage are the same as those when $\Delta=3$. Here pd-EMEE is substantially more efficient than EMEE: for example, when $n=100$, the relative efficiency between pd-EMEE and EMEE for estimating $\beta_{0}$, $\beta_{1}$, and $\beta_{2}$ is $1.450.065^{2}/0.054^{2}=$ 1.45, $1.390.092^{2}/0.078^{2}=$ 1.39, and $1.400.064^{2}/0.054^{2}=$ 1.40, respectively, which roughly translates to 31%, 28%, and 29% savings in sample size. When $\Delta=10$, pd-EMEE2 further substantially improves the efficiency over pd-EMEE. For example, when $n=100$, the relative efficiency between pd-EMEE2 and pd-EMEE for estimating $\beta_{0}$, $\beta_{1}$, and $\beta_{2}$ is $1.040.054^{2}/0.053^{2}=$ 1.04, $1.140.078^{2}/0.073^{2}=$ 1.14, and $1.170.054^{2}/0.050^{2}=$ 1.17, respectively, which roughly translates to another 3%, 12%, and 15% savings in sample size on top of the savings of pd-EMEE. The efficiency improvement when $\Delta=10$ is greater compared to when $\Delta=3$.
@@
Table 2: Simulation result: Performance of pd-EMEE2, pd-EMEE, EMEE, GEE.ind and GEE.exch when estimating $\beta_{0}$ ($\Delta$ = 3 and 10). Based on 1,000 simulation replications.
@@
  • pd-EMEE, the estimator $\hat{\beta}$ proposed in Section 3; pd-EMEE2, the estimator $\tilde{\beta}$ proposed in Section 4; EMEE, the estimator in Qian and others (2021); GEE.ind, generalized estimating equations with independent working correlation structure; GEE.exch, generalized estimating equations with exchangeable working correlation structure; SD, standard deviation; RMSE, root mean squared error; CP, $95\%$ confidence interval coverage probability before (unadj) and after (adj) small-sample correction.
Table 3: Simulation result: Performance of pd-EMEE2, pd-EMEE, EMEE, GEE.ind and GEE.exch when estimating $(\beta_{1},\beta_{2})$ ($\Delta$ = 3 and 10). Based on 1,000 simulation replications.
@@
  • pd-EMEE, the estimator $\hat{\beta}$ proposed in Section 3; pd-EMEE2, the estimator $\tilde{\beta}$ proposed in Section 4; EMEE, the estimator in Qian and others (2021); GEE.ind, generalized estimating equations with independent working correlation structure; GEE.exch, generalized estimating equations with exchangeable working correlation structure; SD, standard deviation; RMSE, root mean squared error; CP, $95\%$ confidence interval coverage probability before (unadj) and after (adj) small-sample correction.
We then investigated how the relative efficiency of pd-EMEE or pd-EMEE2 against EMEE depends on $\Delta$ (proximal outcome window length) and $p_{a}$ (randomization probability) by varying $\Delta$ and $p_{a}$ separately. When $\Delta$ varies from 1 to 10 with $p_{a}=0.2$, the relative efficiency increases from 1 to 1.5 for pd-EMEE and 1 to 1.75 for pd-EMEE2 (left panel of Figure 2). When $p_{a}$ varies from 0.1 to 0.8 with $\Delta=3$, the relative efficiency increases from 1.05 to 1.75 for pd-EMEE and from 1.06 to 2.00 for pd-EMEE2 (right panel of Figure 2). Therefore, a larger $\Delta$ or a larger $p_{a}$ makes the proposed pd-EMEE and pd-EMEE2 even more efficient than the EMEE method. This phenomenon was expected and explained in Section 3.
Figure 2: Simulation result: The relative efficiency (RE) of pd-EMEE (red) or pd-EMEE2 (blue) over EMEE and how it varies with $\Delta$ (proximal outcome window length) and $p_{a}$ (randomization probability). The variances of EMEE, pd-EMEE, and pd-EMEE2 are illustrated by various line types in black.